Find the LCM and HCF of the following pairs of integers and verify that LCM × HCF = product of the two numbers.   336 and 54

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Expressing 336 and 54 as product of its prime factors, we get,

336 = 2 × 2 × 2 × 2 × 7 × 3 × 1
54 =2×3×3×3×1
Therefore,
LCM(336, 54) =24× 33× 7= 3024
AndHCF(336,54) = 2× 3 = 6
Verification
Now, product of 336 and 54 = 336× 54 =18,144


Product of LCM and HCF =3024 ×6 = 18,144

Hence,
LCM × HCF = product of the 336 and 54.

Thanks